OpenStudy (anonymous):

A tennis player throws a ball upward at 15 m/s. What is the velocity after 1.0s. (Please consider that gravity is 9.8m/s^2 going down)

OpenStudy (anonymous):

Look and see what variables you have and dont have. Velocity initial Acceleration Time No velocity final. Use your formula sheet and look to see which formula you can use to figure the Velocity final is.

OpenStudy (anonymous):

does anything happen with the initial velocity since it is going upward? I assumed it would be negative, but I could be wrong?

OpenStudy (anonymous):

if a ball travels 15m/s and is losing speed at 9.8m/s^2, you can just subtract it off. You happen to be subtracting 9.8 from 15. In more general sense. V2m/s =V1m/s - (Gm/s^2)*Ts = V1m/s - Gm/s I wrote that out so you can see what happens to the units. In this case just sub 1 second into T, for time.

OpenStudy (anonymous):

use formula v=u+at v=final velocity u=intial velocity acceleration will b negative...

OpenStudy (anonymous):

Tyvm, got it solved.